# Mott-glass phase induced by long-range correlated disorder in a one-dimensional Bose gas

### v1

@@ Introduction.
Figure 2: Flow diagrams in the space $(K,\sqrt{\lambda_{2}},\sqrt{\delta_{1}})$ for $\sigma=-0.2$ (left) and $\sigma=-0.4$ (right). The black dots show the BKT fixed point $((3-\sigma)/2,0,0)$ controlling the transition between the superfluid phase and the localized phase, as well as the Bose-glass (BG) and Mott-glass (MG) fixed points. (The dashing of some lines is used for improved visibility and has no particular physical meaning.)
@@
(we have dropped the $k$ index in $\Delta_{i}$, $K$, $\theta$ and $z$ for ease of notations), where
@@
Figure 3: Flow trajectories, projected in the plane $(K,\sqrt{\lambda_{2}})$, near the BKT fixed point located at $K_{c}=1.7$ ($\sigma=-0.4$). The initial conditions are defined by $\delta_{1,\Lambda}\sim 0.1$ when $K_{\Lambda}>1.7$, and $\delta_{1,\Lambda}\sim 10^{-6}$ when $K_{\Lambda}<1.7$. The dashed lines show the two separatrix obtained from the standard BKT flow equations derived from (9-13) to leading order in $K_{k}-K_{c}(\sigma)$ and $\sqrt{\lambda_{2,k}}$ [22].
@@ Conclusion.
Thus we believe that long-range disorder correlations appear as the most natural way proposed so far to observe a Mott glass in a one-dimensional Bose gas. Cold atoms are a natural platform for experimental study. Laser speckles allow to control, to some extent, the disorder potential acting on the gas [46, 47], but long-range power-law correlations seem out of reach. On the other hand, by generating the disorder potential from the spatial modulation of a laser beam using a digital micromirror device (DMD), the statistical properties of the disorder potential can be adjusted at will, which could open up the possibility of creating disorder potentials with the desired characteristics (see, e.g., Refs. [48, 49, 50]). Finally we point out that a classical analogue of the Mott glass is expected to exist in heavy-ion-irradiated type-II superconductors with algebraically correlated columnar disorder [22].

### v2

@@ Introduction.
Figure 2: Flow diagrams in the space $(K,\sqrt{\lambda_{2}},\sqrt{\delta_{1}})$ for $\sigma=-0.2$ (left) and $\sigma=-0.4$ (right). The black dots show the BKT fixed point $((3-\sigma)/2,0,0)$ controlling the transition between the superfluid phase and the localized phase, as well as the Bose-glass (BG) and Mott-glass (MG) fixed points. (The dashing of some lines and the different colors are used to improve visibility and have no particular physical meaning.)
@@
(we have dropped the $k$ index in $\Delta_{i}$, $K$, $\theta$, and $z$ for ease of notations), where
@@
Figure 3: Flow trajectories, projected in the plane $(K,\sqrt{\lambda_{2}})$, near the BKT fixed point located at $K_{c}=1.7$ ($\sigma=-0.4$). The initial conditions are defined by $\delta_{1,\Lambda}\sim 0.1$ when $K_{\Lambda}>1.7$, and $\delta_{1,\Lambda}\sim 10^{-6}$ when $K_{\Lambda}<1.7$. The dashed lines show the two separatrixes obtained from the standard BKT flow equations derived from (9-13) to leading order in $K_{k}-K_{c}(\sigma)$ and $\sqrt{\lambda_{2,k}}$ [22]. The different colors of the trajectories are used to improve visibility and have no particular physical significance.
@@ Conclusion.
Thus we believe that long-range disorder correlations appear as the most natural way proposed so far to observe a Mott glass in a one-dimensional Bose gas. Cold atoms are a natural platform for experimental study. Laser speckles allow to control, to some extent, the disorder potential acting on the gas [46, 47], but long-range power-law correlations seem out of reach. On the other hand, by generating the disorder potential from the spatial modulation of a laser beam using a digital micromirror device (DMD), the statistical properties of the disorder potential can be adjusted at will, which could open up the possibility of creating disorder potentials with the desired characteristics (see, e.g., Refs. [48, 49, 50]). Finally we point out that a classical analog of the Mott glass is expected to exist in heavy-ion-irradiated type-II superconductors with algebraically correlated columnar disorder [22].